OpenStudy (anonymous):

using a calculator find tan theta if sec theta=-2.24

OpenStudy (anonymous):

Well \[ \sin(\theta)=-\frac{1}{2.24} \]

OpenStudy (anonymous):



OpenStudy (anonymous):

You can find the top side using Pythagorean theorem.

OpenStudy (anonymous):

Tangent is left side / top side

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \tan(\theta)=-\frac{1}{\sqrt{(2.24)^2-1^2}{}} \]

OpenStudy (anonymous):

I found it out by using the identity "1+tan^2=sec^2".

OpenStudy (anonymous):

Oh, that works.